# The exterior derivative and the mean value equality in $\mathbb{R}^{n}$

### v1

@@ Abstract
This survey revisits classical results in vector calculus and analysis by exploring a generalised perspective on the exterior derivative, interpreting it as a measure of “infinitesimal flux”. This viewpoint leads to a higher-dimensional analogue of the Mean Value Theorem, valid for differential $k$-forms, and provides a natural formulation of Stokes’ theorem that mirrors the exact hypotheses of the Fundamental Theorem of Calculus — without requiring full $C^{1}$ smoothness of the differential form.
@@ 1 Introduction
Let $U\subset\mathbb{R}$ be an open subset and $f:U\to\mathbb{R}$ be a real function; interpreting any closed interval $a,b]\subseteq U$ as an oriented $1$-chain with boundary (see Section 2.2 for definitions), we denote
@@
f $f\text{ is differentiable at }x\in U\iff\exists\lim_{[a,b]\to x}\frac{1}{\text{\rm vol}[a,b]}\int_{\partial[a,b]}$f,
@@
$a,b]\subseteq U,x\in[a,b],\text{\rm vol}[a,b]<\delta\Rightarrow\left|\frac{f(b)-f(a)}{b-a}-f^{\prime}(x)\right|<\varepsilon.$
@@
In the general setting of $k$-forms in dimension $n$, we mimic this lemma and define the exterior derivative of $\omega$ at a point $x$, evaluated on the generators of a typical cube in $\mathbb{R}^{k+1}$, to be a limit of the ‘average’ flux of the $k$-form $\omega$ across the boundary of suitable shrinking $k$-blocks (see Definition 11). Now, in dimension $n>1$, the condition on the right-hand side of such definition still makes sense even if $\omega$ is not continuous at $x$, let alone differentiable, so this notion extends the usual exterior derivative to a larger class of differential forms. Moreover, with this viewpoint, we can prove the following generalisation of MVT as an equality (see Theorem 5):
@@
    $\omega$ is derivable on $\mathring{B}$, i.e. $\omega$ admits an exterior derivative $\text{\rm D}\omega$ in the broader sense of Definition 11, which coincides with the standard $d\omega$ in the differentiable case (see Theorem 8);
@@ 3.2 The infinitesimal flux interpretation
which, by naturality of the pullback, can be rewritten as
@@
$\omega=\sum_{I}\omega_{I}dx^{I}\in\Omega^{k-1}(U),$

### v2

@@ Abstract
This survey revisits classical results in vector calculus and analysis by exploring a generalised perspective on the exterior derivative, interpreting it as a measure of “infinitesimal flux”. This viewpoint leads to a higher-dimensional analogue of the Mean Value Theorem, valid for differential $k$-forms, and provides a natural formulation of Stokes’ theorem that mirrors the exact hypotheses of the Fundamental Theorem of Calculus – without requiring full $C^{1}$ smoothness of the differential form.
@@ 1 Introduction
Let $U\subset\mathbb{R}$ be an open subset and $f:U\to\mathbb{R}$ be a real function. Interpreting any closed interval $a,b]\subseteq U$ as an oriented $1$-chain with boundary (see Section 2.2 for definitions), we denote
@@
$f\text{ is differentiable at }x\in U,\text{ with }f^{\prime}(x)=L\iff\exists\lim_{[a,b]\to x}\frac{1}{\text{\rm vol}[a,b]}\int_{\partial[a,b]}f=L,$
@@
$a,b]\subseteq U,x\in[a,b],\text{\rm vol}[a,b]<\delta\Rightarrow\left|\frac{f(b)-f(a)}{b-a}-L\right|<\varepsilon.$
@@
In the general setting of $k$-forms in dimension $n$, we mimic Lemma 2 and define the exterior derivative $\text{\rm D}\omega_{x}$ of $\omega$ at the point $x$ by requiring that, for every choice of directions $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$, the value of $(\text{\rm D}\omega)_{x}(v_{1},\dots,v_{k})$ be given by the limit of the normalised flux of $\omega$ across the boundary of $C^{1}$-parametrised $k$-blocks shrinking to $x$ (see Definition 11). In dimensions $n>1$ and degree $k>0$, the right-hand side of this definition still makes sense even when $\omega$ is not continuous at $x$, let alone differentiable (see e.g. Example 2). Moreover, we prove that in the differentiable case this flux-based exterior derivative $\text{\rm D}\omega$ agrees with the usual exterior derivative $\text{\rm d}\omega$ (see Theorem 8). Thus, the definition extends the classical operator d to a strictly larger class of differential forms. Moreover, with this viewpoint, we can prove the following generalisation of MVT as an equality (see Theorem 5):
@@
    $\omega$ is derivable on $\mathring{B}$, i.e. $\omega$ admits an exterior derivative $\text{\rm D}\omega$ in the broader sense of Definition 11, which agrees with the usual $\text{\rm d}\omega$ when $\omega$ is differentiable (Theorem 8);
@@ 2.2 Chains, blocks and integration in $\mathbb{R}^{n}$
Remark 3.
Using Definitions 4 and 5, it is easy to check that if $\omega\in\Omega^{k-1}(U)$, where $U\subset\mathbb{R}^{n}$ is an open subset, and $c:B\to U$ is a singular $k$-block of class $C^{1}$, then
$\int_{\partial c}\omega=\int_{\partial B}c^{\ast}\omega,$
provided the (Riemann) integrals exist.
@@ 3.2 The infinitesimal flux interpretation
which can be rewritten as
@@
$\omega=\sum_{I}\omega_{I}\text{\rm d}x^{I}\in\Omega^{k-1}(U),$